OpenStudy (anonymous):

Compare the functions shown below: f(x) = -3x − 5 h(x) = 2 cos(2x − π) + 4

OpenStudy (anonymous):

g(x)

OpenStudy (anonymous):

Which function has the greatest y-intercept? f(x) g(x) h(x) All three functions have the same y-intercept.

OpenStudy (agreene):

So, the y intercept is where x=0 on g(x) we can look at the graph and know it is -3 \(f(0)=-3(0)-5 \rightarrow -5\) \[h(0)=2\cos(0-\pi)+4 \rightarrow-2+4\rightarrow2\] so we have f(0) = -5 g(0) = -3 h(0) = 2

OpenStudy (anonymous):

@agreene Thank you soo much!